The indices of refraction of the core and cladding of an optical fiber are 1.5 and 1.40, respectively. Light is incident at the core at 70%. What is the distance between two successive reflections in the core it its diameter is 0.05mm?
1
Expert's answer
2020-05-18T10:16:55-0400

As per the question,

Refractive index of core "(\mu_1)=1.5"

Refractive index of cladding "(\mu_2)=1.4"

Angle of incidence "= 70^\circ"

Diameter of the core"=0.05mm"

"\Rightarrow L=d\sqrt{(\dfrac{\mu_2}{\mu_1\sin\theta})^2-1}"

"\Rightarrow L=0.05\sqrt{(\dfrac{1.5}{1.4\sin(70)})^2-1}"

"\Rightarrow L =0.05\sqrt{1.27-1}=0.05\sqrt{0.27}=0.025mm"